OpenStudy (anonymous):

Please I am completely lost with this problem it states to add the polynomials / show all work please. (18x^2 - 1/4x - 3/8) + (-1/2x^2 - 1/4x - 1/2)

OpenStudy (anonymous):

\[18x^2-\frac{1}{2}x^2=17.5x^2\] \[-\frac{1}{4}x-\frac{1}{4}x=-\frac{1}{2}x\] \[-\frac{3}{8}-\frac{1}{2}=-\frac{7}{8}\]

OpenStudy (anonymous):

answer is \[17.5x^2-\frac{1}{2}x-\frac{7}{8}\]